how many mega coulombs of positive (or negative) charge are in 1.00 mole of neutral molecular hydrogen gas ?


 

Grade:

3 Answers

pankaj kumar raman
12 years ago

or it will contain 2.00 mol of hydrogen nuclei/electron/proton

1 mol electron/proton has 1farad charge=96500coulomb=9.65×10e-2 MCoulomb

hence 2.00 mol will have 0.193MCoulomb charge

ashutosh aggarwal
12 years ago

In every molecule of H2 there are 2 electrons.So total no. of electrons in 1 mole is = 2*6.022*1023 .

As charge on 1 e is 1.6*10-19 coulombs.

Thus total charge = 2*6.022*1023*1.6*10-19 coulombs=19.2704*104 coulombs=19.2704*10-2 mega coulombs.

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
1 mole H2 = 2.00 mol of hydrogen atom = 2 mol of nuclei/electron/proton
1 mol electron/proton has 1 farad charge = 96500coulomb = 9.65×10-2 M Coulomb
Hence 2.00 mol will have 0.193MCoulomb charge
 
Hope it helps.
Thanks and regards,
Kushagra